find the altitude of a triangle if its base has length 10 and the triangle is equal in are to a parallelogram whose base and altitude have lengths 15 and 8
Answer by Cromlix(4381)

Area of parallelogram = length * width
Area of parallelogram = 15 * 8
Area of parallelogram = 120 units^2
Area of triangle = 1/2 (base * height)
120 units^2 = 1/2 ( 10 * height)
120 * 2 = 10 * height
Height = 240/10
Height of altitude = 24 units.
Hope this helps.
:-)